OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

you mean "sine - squared -x"?

OpenStudy (anonymous):

\[7\sin^2x-14\sin x+2=-5\]

OpenStudy (anonymous):

Yes, that's what I meant.

OpenStudy (anonymous):

7sin^2x-14sinx+2=-5 7sin^x-14sinx+7=0 sin^2x-2sinx+1=0 (sinx-1)^2=0 sinx-1=0 sinx=1=sin\[\frac{ \pi }{ 2 }\] x=\[\frac{ \pi }{ 2 }\]